OpenStudy (anonymous):

a bag contains 6 red marbles, 6 white marbles, and 4 blue marbles find p(red or blue)

OpenStudy (anonymous):

\[P(RED \cup BLUE)\]

OpenStudy (anonymous):

\[P(X\cup Y) = P(X) +P(Y)-P(X \cap Y)\]

OpenStudy (anonymous):

in this case X = RED Y = BLUE

OpenStudy (anonymous):

P(RED) = 6/(6+6+4) = 6/16 P(BLUE) = 4/(6+6+4) = 4/16 P(Drawing a Red and a Blue marble in the same draw) = 0/16 do the addition to find the probability.

OpenStudy (kropot72):

There is a total of 16 marbles. 6 are red marbles and 4 are blue marbles. In a single draw the probability of drawing a red marble is 6/16 and the probability of drawing a blue marble is 4/16. It is not possible to draw both colors in a single draw. Therefore P(red) and P(blue) are said to be mutually exclusive. To find the probability of a red or a blue marble we simply add the two values of probability: \[P(red\ or\ blue)=P(red)+P(blue)=\frac{6}{16}+\frac{4}{16}= you\ can\ calculate\]